# NoisyRollout: Reinforcing Visual Reasoning with Data Augmentation

### v1

@@ 4 Related Work
4.1 Large Vision-Language Models
VLMs have rapidly evolved to understand and reason with both visual and textual information. These models combine visual encoders with large language models to enable comprehension and inference across modalities. Early VLMs like Flamingo (Alayrac et al., 2022) and BLIP-2 (Li et al., 2023) established foundational integration techniques between vision and language components. The LLaVA series (Liu et al., 2023, 2024b; Li et al., 2024b, a) introduced effective visual instruction tuning methodologies that significantly advanced multimodal capabilities. For mathematical reasoning, specialized approaches like Math-LLaVA (Shi et al., 2024) and MAVIS (Zhang et al., 2024b) have employed mathematical visual instruction tuning to enhance VLMs’ abilities to interpret and solve mathematical problems in multimodal contexts.
Advanced VLMs including GPT-4o (Hurst et al., 2024) and Gemini (Gemini Team et al., 2023) have demonstrated unprecedented general visual understanding through massive pretraining. Mixture-of-Experts approaches in DeepSeek-VL2 (Wu et al., 2024b), Uni-MoE (Li et al., 2025), and MoVA (Zong et al., 2024) improved computational efficiency by selectively activating specialized components based on input characteristics. Meanwhile, unified models like SEED-X (Ge et al., 2024), Chameleon (Team, 2024), Show-O (Xie et al., 2024), and Janus series (Wu et al., 2024a; Ma et al., 2024; Chen et al., 2025c) integrated visual understanding and generation capabilities within single architectures. However, most existing VLMs still lack robust visual reasoning capabilities, especially for tasks requiring sophisticated analysis of visual information combined with complex reasoning.
4.2 Reinforcement Learning in LLMs and VLMs
RL has emerged as a key methodology for enhancing the capabilities of LLMs and VLMs. Early research primarily focused on Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), which aligned model outputs with human preferences (Achiam et al., 2023). Recent advancements have further demonstrated that RL-based techniques can significantly enhance reasoning abilities. For instance, DeepSeek-R1 (Guo et al., 2025) utilizes rule-based rewards combined with Group Relative Policy Optimization (GRPO) (Shao et al., 2024), whereas Kimi-1.5 (Kimi Team, 2025) employs a variant of online policy mirror descent, both methods showing notable improvements in reasoning performance. In the multimodal domain, research on leveraging RL to enhance VLMs’ reasoning capabilities remains in early stages. Some approaches explore using generative reward models (Zhang et al., 2025b) to enhance VLMs’ general capability, but these typically require powerful closed-source models for training data generation. Recent work including LMM-R1 (Peng et al., 2025), Vision-R1 (Huang et al., 2025), R1-V (Chen et al., 2025b) and OpenVLThinker (Deng et al., 2025) has applied R1-style reinforcement learning to VLMs in specific subdomains like geometry problems and object counting tasks. MM-Eureka (Meng et al., 2025) further extends large-scale rule-based RL to broader multimodal mathematical reasoning, demonstrating significant performance gains without relying on in-domain training data.
@@ 5 Conclusion
In this paper, we investigate scaling test-time compute in VLMs via RL. We propose NoisyRollout, a simple rollout strategy that introduces rollout diversity through clean and distorted inputs from the visual perspective, enhancing policy exploration during RL training. Without modifying the RL objective, NoisyRollout improves generalization and robustness, achieving state-of-the-art results on multiple benchmarks with minimal training data. In future work, we plan to evaluate NoisyRollout across diverse RL objectives and datasets, and explore adaptive noise scheduling to better accommodate varying training conditions.
Appendix A Appendix
A.1 Additional Ablation Study on the Strategies of Noise Annealing
On the Geometry3K training dataset, we further examine the impact of different noise annealing schedules on NoisyRollout’s performance by comparing our default sigmoid strategy with power ($\alpha_{t}=\alpha_{0}\cdot(1-t/T_{\text{max}})^{p}$, $p=3.0$) and exponential ($\alpha_{t}=\alpha_{0}\cdot\gamma^{t/T_{\text{max}}}$, $\gamma=0.98$) decay functions. As shown in Table 6, all three strategies enable NoisyRollout to outperform the vanilla GRPO baseline. However, the sigmoid schedule consistently achieves the highest performance, reaching an average score of $58.6$%, compared to $57.1$% for power decay and $57.0$% for exponential decay. The superior performance of the sigmoid schedule likely stems from its characteristic “slow-fast-slow” decay profile. This shape maintains higher noise levels early on to encourage exploration, rapidly decreases noise during the main convergence phase. This contrasts with the monotonic nature of power and exponential decays, suggesting that the specific profile of noise reduction significantly influences the balance between exploration and stability, thereby impacting the final model effectiveness.
Table 6: Ablation study on the strategies of noise annealing. “Avg.” denotes the average accuracy across all six visual reasoning benchmarks. Best value per column is bold, second best is underlined.
@@ A.1 Additional Ablation Study on the Strategies of Noise Annealing
+ NoisyRollout (Power) 52.2 52.1 26.4 72.0 68.7 71.0 57.1
+ NoisyRollout (Exponential) 51.9 52.6 27.7 70.5 70.1 69.1 57.0
+ NoisyRollout (Sigmoid) 54.9 53.2 28.5 72.6 69.6 72.1 58.6
A.2 Sensitivity to Training Data Variations
To evaluate the generalization of our findings, we first examine the performance consistency of NoisyRollout when using different random seeds for data sampling during training on the Geometry3K dataset. We compare these results against vanilla GRPO trained with the same seeds to verify that the improvements offered by NoisyRollout are robust and not merely an artifact of a specific data ordering or subset selection (Table 7). We then assess the generalizability of NoisyRollout by applying it to a different training dataset, K12. This experiment tests whether the benefits of our noisy training strategy extend beyond the Geometry3K domain. Furthermore, within this K12 training setup, we investigate the model’s sensitivity to varying initial noise steps to understand how the optimal noise level might depend on the dataset characteristics (Table 8).
@@ A.2 Sensitivity to Training Data Variations
Table 7: Performance comparision of NoisyRollout and vanilla GRPO across different data seeds. “Avg.” denotes the average accuracy across all six benchmarks.
Method Seed Geo3K MathVerse MathVision MathVista WeMath HalluBench Avg.
GRPO 1 52.0 50.1 27.6 69.4 66.6 69.4 55.9
NoisyRollout 54.9 53.2 28.5 72.6 69.6 72.1 58.6
GRPO 2 51.9 51.5 27.8 70.7 68.1 69.4 56.6
NoisyRollout 54.1 53.6 26.6 72.9 69.7 71.4 58.1
GRPO 42 51.9 50.5 27.2 71.3 68.6 68.6 56.4
NoisyRollout 54.1 52.7 27.3 72.2 69.6 70.3 57.7
@@
Table 8: Performance of NoisyRollout on the K12 dataset across different distortion levels. “Avg.” denotes the average accuracy across all six benchmarks
@@
Method Noise Step Geo3K MathVerse MathVision MathVista WeMath HalluBench Avg.
Qwen2.5-VL-7B-Instruct N/A 39.4 46.2 25.3 67.5 63.0 64.6 51.0
NoisyRollout 00 39.6 50.7 28.5 71.7 68.6 69.8 54.8
$300$ 39.9 51.4 28.4 72.5 69.5 70.5 55.4
$400$ 39.8 50.3 28.2 71.8 70.3 70.8 55.2
$450$ 43.9 52.8 28.9 72.9 71.9 70.7 56.9
$500$ 41.4 50.2 27.4 73.2 71.4 71.3 55.8
A.3 Evaluating the Perception Quality of Reasoning Traces
@@
To further evaluate the perception quality of models trained with NoisyRollout and vanilla GRPO during reasoning, we perform a paired comparison using a strong vision-language model [^7] Specifically, we use Gemini-Flash-2.0-001 in this experiment.. We sample 300 reasoning traces from the evaluation logs of the models performing visual reasoning on the MathVerse and MathVista benchmarks, forming paired comparisons between NoisyRollout and vanilla GRPO. To isolate visual perception, we extract only the visual components from each reasoning trace using a specialized prompt, removing any influence from mathematical reasoning or final answers. To reduce potential position bias in the comparisons, each pair of traces is evaluated twice: once with the NoisyRollout trace shown first and the vanilla GRPO trace second, and once with the order reversed. We combine the results using the Bradley-Terry model to compute win rates. This methodology offers a reliable measure focused specifically on visual perception quality during reasoning. The results are presented in Figure 3 (the $8$th subfigure). The extraction and evaluation prompts are shown above.
A.4 Case Study

### v3

@@ 4 Related Work
VLMs have rapidly advanced through integrating vision encoders [54, 80] with large language models [2, 32, 36, 37, 22, 17], with specialized efforts in reasoning tasks [58, 83]. Reinforcement learning training in LLMs and VLMs, initially employed for alignment via human feedback (RLHF) [49, 1, 77], has evolved to incorporate rule-based rewards and advanced optimization methods like GRPO [56], as exemplified by DeepSeek-R1 [20] and Kimi-1.5 [26]. Emerging RL approaches in multimodal domains include LMM-R1 [51], Vision-R1 [21], R1-V [7], OpenVLThinker [13], and MM-Eureka [48], which extend RL to visual reasoning tasks. However, existing studies on training VLMs via RL have not adequately explored techniques that can enhance the explorative capabilities of models. Our method addresses this gap by proposing a data augmentation technique with visual-oriented inductive biases. A detailed discussion of related work is deferred to the Appendix E due to space limit.
@@ 5 Conclusion
In this paper, we investigate scaling test-time compute in VLMs via RL. We introduce NoisyRollout, a simple yet effective data augmentation technique that promotes diversity by mixing trajectories from both clean and distorted inputs with vision-oriented inductive biases. This approach enhances policy exploration during RL training without incurring additional training costs. Empirically, NoisyRollout demonstrates improved generalization and robustness, achieving state-of-the-art performance across multiple visual reasoning and perception benchmarks with high sample efficiency.
Appendix A Additional Ablation Studies
Noise annealing strategy. On the Geometry3K training dataset, we further examine the impact of different noise annealing schedules on NoisyRollout’s performance by comparing our default sigmoid strategy with power ($\alpha_{t}=\alpha_{0}\cdot(1-t/t_{\text{max}})^{p}$, $p=3.0$) and exponential ($\alpha_{t}=\alpha_{0}\cdot\gamma^{t/t_{\text{max}}}$, $\gamma=0.98$) decay functions. As shown in Table 7, all three strategies enable NoisyRollout to outperform the vanilla GRPO baseline on benchmarks. Among them, the sigmoid schedule achieves the highest average score ($58.5$%), surpassing both power and exponential decays ($57.1$% and $57.0$%). The superior performance of the sigmoid schedule likely results from its characteristic “slow-fast-slow” decay, which balances exploration and stability more effectively by maintaining sufficient early exploration and promoting rapid convergence afterward.
Table 7: Ablation study on the strategies of noise annealing. “Avg.” denotes the average accuracy across the six benchmarks. Best value per column is bold, second best is underlined.
@@ Appendix A Additional Ablation Studies
+ NoisyRollout (Pow.) 52.2 52.1 26.4 72.0 68.7 71.0 57.1
+ NoisyRollout (Exp.) 51.9 52.6 27.7 70.5 70.1 69.1 57.0
+ NoisyRollout (Sig.) 54.9 53.2 28.5 72.6 69.6 72.1 58.5
Total rollout number. We analyze the impact of total rollout number by comparing vanilla GRPO and NoisyRollout under varying rollout budgets. As shown in Table 8, increasing the number of rollouts in vanilla GRPO from $n=8$ to $n=16$ improves in-domain performance (from $49.6\%$ to $54.7\%$), but only marginally benefits out-of-domain generalization (from $56.8\%$ to $57.5\%$). NoisyRollout consistently outperforms vanilla GRPO even when the total number of rollouts is held constant. Notably, NoisyRollout with $n_{1}=n_{2}=6$ (total $12$) achieves both higher in-domain ($54.9\%$) and out-of-domain ($59.2\%$) accuracy than vanilla GRPO with $16$ rollouts.
@@
Table 8: Comparision of NoisyRollout and GRPO on Qwen2.5-VL-7B-Instruct across different rollout configurations. “OOD Avg.” denotes the average accuracy across all five out-of-domain benchmarks.
Rollout Geo3K MathVerse MathVision MathVista WeMath HalluBench OOD Avg.
$n=8$ 49.6 50.2 27.3 68.8 68.1 69.7 56.8
$n=12$ 52.0 50.8 27.3 70.5 67.4 69.8 57.2
$n=16$ 54.7 51.3 27.5 71.4 68.0 69.3 57.5
$n=20$ 53.6 50.8 27.2 70.0 68.5 69.9 57.3
$n_{1}=n_{2}=4$ 51.7 51.8 27.4 72.6 69.5 70.3 58.3
$n_{1}=n_{2}=6$ 54.9 53.2 28.5 72.6 69.6 72.1 59.2
$n_{1}=n_{2}=8$ 54.7 52.6 28.7 72.0 69.1 72.0 58.9
Figure 9: Comparison of different image augmentation strategies.
Image data augmentation. We explore a range of image distortions beyond Gaussian noise, including cropping and rotation. However, these augmentations often led to critical information loss, resulting in rollouts with consistently zero rewards. This caused unreliable policy gradient estimates and ultimately led to training instability and divergence (Figure 9). We also experiment with randomized augmentation, where one distortion (cropping, rotation, or Gaussian noise) is randomly selected at each training step. However, this approach fails to improve stability. In contrast, Gaussian noise alone preserved essential visual information while introducing moderate perturbations. We find that Gaussian noise serves as an effective regularizer during training: rollouts generated with this type of augmentation closely resemble those encountered during test-time on benchmark tasks, thereby injecting useful variability into the inputs.
Table 9: Comparision of NoisyRollout and GRPO on Qwen2.5-VL-7B-Instruct across data seeds.
Method Seed Geo3K MathVerse MathVision MathVista WeMath HalluBench OOD Avg.
GRPO 1 52.0 50.8 27.3 70.5 67.4 69.8 57.2
NoisyRollout 54.9 53.2 28.5 72.6 69.6 72.1 59.2
GRPO 2 51.9 51.5 27.8 70.7 68.1 69.4 57.5
NoisyRollout 54.1 53.6 26.6 72.9 69.7 71.4 58.8
GRPO 3 51.1 50.8 27.1 69.7 67.8 69.5 57.0
NoisyRollout 53.7 52.4 27.8 72.3 70.5 70.9 58.8
GRPO 42 51.9 50.5 27.2 71.3 68.6 68.6 57.2
NoisyRollout 54.1 52.7 27.3 72.2 69.6 70.3 58.4
Data seed. To evaluate the robustness of our approach to data sampling variations, we examine the performance consistency of NoisyRollout when using different random seeds for data sampling during training on the Geometry3K dataset. We compare these results against vanilla GRPO trained with the same seeds to verify that the improvements offered by NoisyRollout are consistent across different data orderings and not merely an artifact of a specific seed (Table 9).
@@
Table 10: Performance of NoisyRollout on the MMK12 dataset ($2.1$K) across different noise steps.
@@
Method Initial Noise Step MathVerse MathVision MathVista WeMath HalluBench OOD Avg.
GRPO 00 50.7 28.5 71.7 68.6 69.8 57.9
NoisyRollout $300$ 51.4 28.4 72.5 69.5 70.5 58.5
$400$ 50.3 28.2 71.8 70.3 70.8 58.3
$450$ 52.8 28.9 72.9 71.9 70.7 59.4
$500$ 50.2 27.4 73.2 71.4 71.3 58.7
Initial noise step (MMK12). Based on our experience, the initial noise step $\alpha_{0}$ is a critical hyperparameter. In addition to the Geometry3K dataset, we also evaluate the impact of different initial noise steps on the MMK12 dataset using Qwen2.5-VL-7B-Instruct. Table 10 shows NoisyRollout outperforms standard GRPO on MMK12 when configured with an appropriate initial noise step ($\alpha_{0}=450$).
Proportion of noisy rollouts. Table 11 demonstrates that incorporating noisy rollouts during training significantly enhances model performance across both reasoning and perception benchmarks. A balanced $/50$ 50 distribution of clean and noisy rollouts achieves optimal results ($58.5$% average accuracy), outperforming both the no-noise baseline ($56.3$%) and higher noise proportions. This finding aligns with our earlier observations about the effectiveness of noise as a regularizer, confirming that the ideal approach combines clean rollouts for exploitation of current state with noisy rollouts for exploration, rather than relying exclusively on either strategy.
Table 11: Comparison across different proportions of noisy rollouts $n_{2}/(n_{1}+n_{2})$ during rollout collection, with a fixed total of $12$ rollouts.
Proportion Geo3K MathVerse MathVision MathVista WeMath HalluBench Avg.
$0/12$ 52.0 50.8 27.3 70.5 67.4 69.8 56.3
$3/12$ 51.1 52.8 28.5 71.8 69.8 70.6 57.4
$6/12$ 54.9 53.2 28.5 72.6 69.6 72.1 58.5
$9/12$ 54.2 52.1 27.3 72.3 71.1 72.3 57.9
$/12$ 12 53.1 52.8 28.5 72.3 69.1 71.1 57.8
Appendix B Unsuccessful Attempts
Optimizing noisy and clean trajectories on corresponding inputs. We explored an alternative design in which policy updates were conditioned on the same input used to generate each rollout—i.e., using clean inputs for clean rollouts and distorted inputs for noisy rollouts. However, this approach did not yield meaningful improvements over the original GRPO baseline. We hypothesize that this design reduces the benefits of group-based advantage estimation. Specifically, by decoupling clean and noisy rollouts during optimization, the method effectively degrades into a form of sample-level data augmentation. This fragmentation weakens the shared reward signal across rollouts, thereby diminishing the informativeness of group-level statistics such as the normalized advantage. As a result, the exploration benefits introduced by noisy rollouts are not fully leveraged during policy updates.
In contrast, our proposed approach treats clean and noisy rollouts as a unified group for advantage calculation, while anchoring all policy optimization to the clean inputs. This design retains the distributional diversity introduced by noise, but preserves a consistent input distribution for policy updates—striking a balance between exploration and stable learning.
Reward penalty on noisy subgroup. We experimented with applying explicit reward penalties (e.g., $-0.1$) to all noisy rollouts, aiming to encourage the model to better capture contrastive learning signals. However, this approach quickly led to training divergence. Rather than improving its core reasoning and perception abilities, the policy model learned to distinguish between clean and noisy rollouts. As a result, the noisy rollouts rapidly became highly “off-policy”, since the model could easily identify. This distributional mismatch destabilized training and undermined the learning.
Appendix C Training Procedure
To better illustrate the workflow of our method, a simplified overview is provided in Algorithm 1.
Algorithm 1 NoisyRollout: Noisy Reinforcement Fine-Tuning
1:Input: Current policy $\pi_{\theta}$, old policy $\pi_{\theta_{old}}$, dataset $p_{\mathcal{D}}$, training steps $t_{\text{max}}$, clean rollout number $n_{1}$, noisy rollout number $n_{2}$, clip parameter $\epsilon$, initial noise strength $\alpha_{0}$, noise scheduler $\eta(\cdot)$, noise transformation function $T(\cdot)$
2:for $t=1$ to $t_{\text{max}}$ do
3:     Sample batch $(I,\mathbf{q})\sim p_{\mathcal{D}}$
5:     Generate distorted images $\tilde{I}=T_{\alpha_{t}}(I)$
9:     Compute advantages $\hat{A}_{i}=\frac{r_{i}-\mathrm{mean}(\mathbf{r})}{\mathrm{std}(\mathbf{r})}$, where $\mathbf{r}=\{r_{i}\}_{i=1}^{n_{1}+n_{2}}$
10:     Update policy using:
11:     $\mathcal{J}(\theta)=\mathbb{E}\Big{[}\frac{1}{n_{1}+n_{2}}\sum_{i=1}^{n_{1}+n_{2}}\min\Big{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o}_{i}\mid I,\mathbf{q})}\hat{A}_{i},\mathrm{clip}\big{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o}_{i}\mid I,\mathbf{q})},1-\epsilon,1+\epsilon\big{)}\hat{A}_{i}\Big{)}\Big{]}$
12:     $\theta\leftarrow\theta-\nabla_{\theta}\mathcal{J}(\theta)$triangleright Update conditioned on clean images only
13:     $\theta_{\text{old}}\leftarrow\theta$triangleright Update old policy parameters
14:end for
Appendix D Evaluating the Perception Quality of Reasoning Traces
@@
To further evaluate the perception quality of models trained with NoisyRollout and vanilla GRPO during reasoning, we perform a paired comparison using a strong VLM. [^8] Specifically, we use Gemini-Flash-2.0-001 in this experiment. We sample 300 reasoning traces from the evaluation logs of the models performing visual reasoning on the MathVerse and MathVista benchmarks, forming paired comparisons between NoisyRollout and vanilla GRPO.
To isolate visual perception, we extract only the visual components from each reasoning trace using a specialized prompt, removing any influence from mathematical reasoning or final answers. To reduce potential position bias in the comparisons, each pair of traces is evaluated twice: once with the NoisyRollout trace shown first and the vanilla GRPO trace second, and once with the order reversed. We combine the results using the Bradley-Terry model to compute win rates. This methodology offers a reliable measure focused specifically on visual perception quality during reasoning. The results are presented in Figure 3 (the $8$th subfigure). The extraction and evaluation prompts are shown above.
Appendix E Detailed Related Work
Large Vision-Language Models. VLMs have rapidly evolved to understand and reason with both visual and textual information [61]. These models combine visual encoders with large language models to enable comprehension and inference across modalities. Early VLMs like Flamingo [2] and BLIP-2 [32] established foundational integration techniques between vision and language components. The LLaVA series [36, 37, 31, 30] introduced effective visual instruction tuning methodologies that significantly advanced multimodal capabilities. For mathematical reasoning, specialized approaches [58, 83] have employed mathematical visual instruction tuning to enhance VLMs’ abilities to interpret and solve mathematical problems in multimodal contexts.
Advanced VLMs including GPT-4o [22] and Gemini [17] have demonstrated unprecedented general visual understanding through massive pretraining. Mixture-of-Experts approaches in DeepSeek-VL-2 [70], Uni-MoE [34], and MoVA [87] improved computational efficiency by selectively activating specialized components based on input characteristics. Meanwhile, unified models like SEED-X [16], Chameleon [60], Show-o [71], and Janus series [69, 47, 8] integrated visual understanding and generation capabilities within single architectures. However, most existing VLMs still lack robust visual reasoning capabilities [14], especially for tasks requiring sophisticated analysis of visual information combined with complex reasoning [38, 65].
Reinforcement Learning-Enhanced Visual Reasoning. RL has emerged as a key methodology for enhancing the capabilities of LLMs and VLMs. Early research primarily focused on Reinforcement Learning from Human Feedback (RLHF) [49], which aligned model outputs with human preferences [1]. Recent advancements have further demonstrated that RL-based techniques can significantly enhance reasoning abilities. For instance, DeepSeek-R1 [20] utilizes rule-based rewards combined with Group Relative Policy Optimization (GRPO) [56], whereas Kimi-1.5 [26] employs a variant of online policy mirror descent, both methods showing notable improvements in reasoning performance.
In the multimodal domain, research on leveraging RL to enhance VLMs’ reasoning capabilities remains in early stages. Some approaches explore using generative reward models [84, 66] to enhance VLMs’ general capability, but these typically require powerful closed-source models for training data generation. Recent work including LMM-R1 [51], Vision-R1 [21], R1-V [7] and OpenVLThinker [13] has applied R1-type RL to VLMs in diverse specific subdomains like geometry problems and object counting tasks [50, 10, 12, 33, 63, 24]. Further more, pilot studies [48, 46] further extend large-scale rule-based RL to broader multimodal mathematical reasoning, demonstrating significant performance gains without relying on in-domain training data.
Appendix F Broader Impacts
Our proposed method, NoisyRollout, introduces a simple yet powerful data augmentation approach designed to improve visual reasoning and perceptual robustness in VLMs. Given its effectiveness, especially noted through strong out-of-domain performance and high sample efficiency, this approach has broad applicability within resource-constrained training scenarios. This is particularly beneficial in domains where acquiring or annotating large-scale datasets is costly or practically challenging, such as medical imaging [29], robotic perception [59], and assistive technologies [15].
Moreover, by enhancing model robustness to visual conditions, our method can also facilitate safer and more reliable deployment of VLMs in real-world applications, potentially leading to more trustworthy human-AI interactions. Furthermore, as our method involves relatively simple augmentation steps without additional computational overhead or complex training protocols, along with strong performance on scaling experiments as shown in Table 2 and Figure 6, it is suitable for integration into existing large-scale training pipelines, supporting broader adoption in both academia and industry.
Appendix G Limitations
Despite easy-to-adopt designs and promising empirical results, our study has several limitations. First, due to computational constraints, our experiments are limited in scale: we primarily explore model sizes up to $32$B parameters and training dataset scales in the order of a few thousand samples. Future work should validate and extend our findings using significantly larger-scale training scenarios—such as models with $72$B parameters or training datasets in the range of hundreds of thousands of sample. Second, NoisyRollout is applied during the RL fine-tuning phase of an already pre-trained VLM. A more fundamental, but vastly more complex, direction would be to explore how the principles of NoisyRollout (i.e., learning from noisy signals in RL) could be integrated into the large-scale pre-training phase of the VLM itself. Finally, while empirically effective, our study lacks a formal theoretical analysis of how NoisyRollout, with its specific hybrid trajectory mixing and noise annealing, affects the exploration-exploitation trade-off and the convergence properties of the RL algorithm. It’s unclear if the introduced noise guarantees broader state-space coverage in a principled way or if certain noise characteristics could inadvertently hinder convergence.
Appendix H Case Study